OpenStudy (anonymous):

solve for x -2(x + 3) = -2(x + 1) - 4

OpenStudy (anonymous):

Expand the expression on both sides.

OpenStudy (anonymous):

Then, combine all the x's on one side, the constants on another side. You will be able to solve for x then.

OpenStudy (mrnood):

@kaylalynn44 this is showing as posted 1 month ago - are you there online?

OpenStudy (anonymous):

yes idk y it says that. it must be a glitch

OpenStudy (mrnood):

ok - on the left of the equagtion there is a term -2(x+3) Can you expand that? on the right side there is a term -2(x+1) can you expand that?

OpenStudy (anonymous):

Do you mean use the distributive property?

OpenStudy (mrnood):

yes

OpenStudy (anonymous):

x -2x + 2 * 3 = -2x +1x - 4

OpenStudy (mrnood):

a good try - but not quite correct when you distribute you have to remember to distribute the sign (i.e. the -) too

OpenStudy (anonymous):

x -2x * 3 +2 = -2x * -2 + 1 -4 ??

OpenStudy (mrnood):

whoa! take it steady -2(x+3) is the same as -2(x) + -2(3) try both sides again

OpenStudy (anonymous):

x -2x + -2 * 3 = -2x + -2 * 1 -4

OpenStudy (mrnood):

ok good that is correct - but it needs some brackets to make it clearer.... x -2x + -2 * 3 = -2x + -2 * 1 -4 x -2x + (-2 * 3) = -2x + (-2 * 1) -4 so write that out with the numbers completed

OpenStudy (anonymous):

x -2x + (-2 * 3) = (-2 * 1) -4 x -2x + -6 = -2 -4

OpenStudy (mrnood):

you missed out the -2x from the right side so we have x-2x-6 =-2x-6 ok? now add 2x to each side

OpenStudy (anonymous):

x - 6 = -6?

OpenStudy (anonymous):

then add 6 to both sides?

OpenStudy (anonymous):

x = 0?

OpenStudy (mrnood):

good now add 6 to each side...

OpenStudy (anonymous):

x = 0

OpenStudy (mrnood):

yes check that back in the original equation

OpenStudy (mrnood):

the original (when x=0) becomes -6=-6 so x=0 is true solution

OpenStudy (anonymous):

0 -2 (0 + 3) = -2 (0 +1) -4 -2 * 3 = -2 +1 * -4 -6 = 2 * -4 -6 = -6

OpenStudy (anonymous):

Thank you so so so much. I think i understand it better. Thank you for your time!

OpenStudy (mrnood):

yw